# Disentanglement-induced multistability

### v2

@@ Abstract
The time evolution of the Bloch vector $\mathbf{k}=\left(\left\langle\sigma_{x}\right\rangle,\left\langle\sigma_{y}\right\rangle,\left\langle\sigma_{z}\right\rangle\right)$ is shown in Fig. 5. As is demonstrated by the plot in Fig. 5(a), in the absence of disentanglement (i.e. $\gamma_{\mathrm{D}}=0$), the steady state is a fixed point. By turning on disentanglement (all other parameters are kept unchanged), the steady state becomes a periodic limit cycle, as is demonstrated by the plot in Fig. 5(b). Due to the similarity between the Hamiltonians (3) and (6), the underlying mechanism responsible for the instability seen in Fig. 5, is similar to the one seen in Fig. 1.
Figure 5: Parallel pumping. While disentanglement is inactive (i.e. $\gamma_{\mathrm{D}}=0$) in (a), the dimensionless disentanglement rate is $\hbar\gamma_{\mathrm{D}}\beta^{-1}/\mathcal{B}^{2}=100$ in (b). For both cases $\hbar\gamma_{\mathrm{H}}/\mathcal{B}=5$ and $\mathcal{J}/\mathcal{B}=1.05$. The green cross symbols represent initial values of the Bloch vector.
Discussion - The current study explores multistability generated by the modified master equation given by Eq. (1). The proposed modified master equation (1) can be constructed for any physical system whose Hilbert space has finite dimensionality. Any candidate master equation modification has to satisfy some legitimizing properties. For the master equation given by Eq. (1), the condition $\mathrm{d}\operatorname{Tr}\rho/\mathrm{d}t=0$ holds provided that $\operatorname{Tr}\rho=1$ (i.e. $\rho$ is normalized), and $\mathrm{d}\operatorname{Tr}\rho^{2}/\mathrm{d}t=0$, provided that $\rho^{2}=\rho$ (i.e. $\rho$ represents a pure state). The first property guaranties norm conservation, whereas positive-semi-definiteness of $\rho$ is ensured by the second property.

### v3

@@ Abstract
The time evolution of the Bloch vector $\mathbf{k}=\left(\left\langle\sigma_{x}\right\rangle,\left\langle\sigma_{y}\right\rangle,\left\langle\sigma_{z}\right\rangle\right)$ is shown in Fig. 5(c) and (d). As is demonstrated by the plot in Fig. 5(c), in the absence of disentanglement (i.e. $\gamma_{\mathrm{D}}=0$), the steady state is a fixed point. By turning on disentanglement (all other parameters are kept unchanged), the steady state becomes a periodic limit cycle, as is demonstrated by the plot in Fig. 5(d). The dependency on the ratio $\mathcal{J}/\mathcal{B}$ of the magnetization $\left\langle\sigma_{x}\right\rangle$ and the two-spin entanglement $\tau$ is shown in Fig. 5 (a) and (b), respectively. Due to the similarity between the Hamiltonians (3) and (6), the underlying mechanism responsible for the instability seen in Fig. 5, is similar to the one seen in Fig. 1.
Figure 5: Parallel pumping. The (a) magnetization $\left\langle\sigma_{x}\right\rangle$ and (b) two-spin entanglement $\tau$ in steady state are plotted as a function of the ratio $\mathcal{J}/\mathcal{B}$. Assumed parameters’ values are $\hbar\gamma_{\mathrm{H}}/\mathcal{B=}5$ and $\hbar\gamma_{\mathrm{D}}\beta^{-1}/\mathcal{B}^{2}=100$. The green cross symbols in (c) and (d) represent initial values of the Bloch vector. While disentanglement is inactive (i.e. $\gamma_{\mathrm{D}}=0$) in (c), the dimensionless disentanglement rate is $\hbar\gamma_{\mathrm{D}}\beta^{-1}/\mathcal{B}^{2}=100$ in (d). For both (c) and (d) $\mathcal{J}/\mathcal{B}=1.05$.
Discussion - The current study explores multistability generated by the modified master equation given by Eq. (1). The proposed modified master equation (1) can be constructed for any physical system whose Hilbert space has finite dimensionality. Any candidate master equation modification has to satisfy some legitimizing properties. For the master equation given by Eq. (1), the condition $\mathrm{d}\operatorname{Tr}\rho/\mathrm{d}t=0$ holds provided that $\operatorname{Tr}\rho=1$ (i.e. $\rho$ is normalized), and $\mathrm{d}\operatorname{Tr}\rho^{2}/\mathrm{d}t=0$, provided that $\rho^{2}=\rho$ (i.e. $\rho$ represents a pure state). The first property guaranties norm conservation, whereas positive-semi-definiteness of $\rho$ is ensured by the second property, together with the relation $\mathrm{d\log}\left(\det\rho\right)/\mathrm{d}t=-2\operatorname{Tr}\left(\Theta-\left\langle\Theta\right\rangle\right)$ [see Eq. (2.196) of Ref. Buks_QMLN ].